OpenStudy (anonymous):

Find the slope of the line whose equation is 7y = 3x + 1. a.7/3 b.1/7 c.3/7

OpenStudy (anonymous):

if 7y=3x+1 y=3x+1/7

OpenStudy (shamil98):

Divide everything by 7. to get your slope-intercept form of y = mx + b. where m = slope

OpenStudy (aggie12341):

y=3/7x+1/7